# Extended BMS representations and strings

### v1

@@ 1 Introduction
The works of McCarthy [13, 14,15,16,17] provided a mathematical classification of the unitary irreducible representations of the BMS group in four dimensions in the absence of the super rotations. A similar analysis was performed for the BMS group in three dimensions [18,19,20] including the super rotations. More recently the results of McCarthy were further discussed [21,22] in relation to the infrared structure of the S-matrix. Despite these results, this mathematical machinery has not yet found concrete applications in the context of scattering theory and the role of super rotations in the representation theory remains unexplored.
@@ 2 Irreducible representations of $BMS_{3}$
$J_{m},J_{n}]=i(m-n)J_{m+n}+i{c_{1}\over 12}m(m^{2}-1)\delta_{n+m,0},\quad$
$J_{m},P_{n}]=i(m-n)P_{m+n}+i{c_{2}\over 12}m(m^{2}-1)\delta_{n+m,0},\quad[P_{m},P_{n}]=0$ (2.0.1)
where $J_{n}$ and $P_{n}$, $n=0,\pm 1,\pm 2,\ldots$ are the super translations and the super rotations respectively. Also $c_{1}$ and $c_{2}$ are the possible central extensions ($c_{1}=0$ and $c_{2}={3\over G}$ for three dimensional gravity in asymptotically flat spacetime).
@@
$\delta p_{n}=-i\sum_{m}\Lambda_{-m}(-n+m)p_{n+m}$ (2.0.5)
@@ 2.1 Massive $BMS_{3}$ representations
We will refer to this as the massive irreducible representation of $BMS_{3}$ even though the ${\bf P}_{\mu}{\bf P}^{\mu}$ Casimir of the Poincaré group does not generalise to the $BMS_{3}$ group in the sense that it is not a $BMS_{3}$ invariant. Similarly in the next section we will refer to the massless $BMS_{3}$ irreducible representations as the one that corresponds to the same choice of momenta as for the massless irreducible representation of the Poincaré group.
@@
We recall that $\varphi_{0}=0$. This irreducible representation is in the list given in [18].
@@
where we have used equation (2.1.4) and $\ldots$ means terms of order $P^{4}$ and higher. At very lowest order $p_{0}=m+\ldots$ and so at lowest order $C(p)=-p_{0}^{2}+\sum_{n,n\not=0}p_{n}p_{-n}+m^{2}=0$. Thus the super momenta of the states obey the constraint $C(p_{n})=0$ and hence the wavefunction in momentum space depends on an infinite number of variables.
@@
Thus $X(z)$ transforms under diffeomorphisms of $z$ as a weighted scalar which is consistent with interpreting $z$ as a parameter labelling a one dimensional object. We can then write the wavefunction in x-space as $\Psi(X(z),a)$. It follows that the massive irreducible representation of $BMS_{3}$ should be viewed as an object of dimension one, that is, a string which is parameterised by $z$. We will discuss this in much greater detail in section (3.2).
We now find the massive irreducible representation when the central charge $c_{2}\not=0$. In this case the momenta vary under the super rotations as
$\delta p_{n}=-i\sum_{m}\Lambda_{-m}(-n+m)p_{n+m}+{ic_{2}\over 12}n(n^{2}-1)\Lambda_{n}\delta_{n+m,0}$ (2.1.15)
Taking into account that the rest frame momentum is given by $p_{n}^{(0)}=m\delta_{n,0}$ we find that
${\delta p_{n}^{(0)}\over m}=ni(2+{c_{2}\over 12}(n^{2}-1))\Lambda_{n}$ (2.1.16)
In the generic case, the isotropy group is the same as above. However, if the central charge is such that $n=\pm\sqrt{1-24/c_{2}}$ are integers, then the isotropy group is ${\cal H}_{mass}^{c}=\{J_{0},J_{n},J_{-n}\}$ which is the algebra of SO(3). These irreducible representations are also in the list of [18]. We note that if $c_{2}$ is positive then there is only one solution, that is, $c_{2}=24$ with $n=0$ but this does not lead to a larger isotropy group. There are only a limited number of solutions if $c_{2}$ is negative, but it is expected in physical applications that it is positive.
The general state in the irreducible representation is given by equation (2.1.3) and its momenta by (2.1.4) provided we also set $\varphi_{n}=0=\varphi_{-n}$. One can then construct the wave function in x space in the same way as above. The situation is a bit like that for supersymmetry where for certain values of the central charge the irreducible representations of super symmetry are smaller.
@@ 2.2 Massless extended $BMS_{3}$ representations
$\delta p_{n}^{(0)}=0=-{i\over\sqrt{2}}(2n\Lambda_{n}+(1-2n)\Lambda_{n-1}-(1+2n)\Lambda_{1+n})$ (2.2.2)
@@
$\quad 4\Lambda_{2}-3\Lambda_{1}-5\Lambda_{3}=0,\quad\quad 4\Lambda_{-2}-3\Lambda_{-1}-5\Lambda_{-3}=0,\quad\ldots$ (2.2.3)
In the BMS3 case we find that equations (2.2.3) can be solved in terms of two real parameters $a$ and $b$ as follows
$\Lambda_{1}=a=\Lambda_{-1},\quad\Lambda_{0}=2b,\quad\Lambda_{2}={2\over 3}(a-b)=\Lambda_{-2},\quad\Lambda_{3}=-{(a+8b)\over 15}=\Lambda_{-3},\quad\ldots$ (2.2.4)
As a result the istropy group for the massless extended $BMS_{3}$ is ${\cal H}_{m=0}^{3}=\{H_{1},H_{2}\}$ where
$H_{1}=J_{0}-{1\over 3}(J_{2}+J_{-2})-{4\over 15}(J_{3}+J_{-3})+\ldots,\quad$
$H_{2}=(J_{1}+J_{-1})+{2\over 3}(J_{2}+J_{-2})-{1\over 15}(J_{3}+J_{-3})+\ldots$ (2.2.5)
@@
$\delta p(z)=i\Big(\Lambda(z)z{d\over dz}p(z)+2p(z)z{d\over dz}\Lambda(z)\Big)$ (2.2.7)
@@
$\delta p^{(0)}(z)=0=i\Big(\Lambda(z)z{d\over dz}p^{(0)}(z)+2zp^{(0)}(z){d\over dz}\Lambda(z)\Big)$ (2.2.8)
@@
${d\over dz}((\Lambda(z)^{2}p^{(0)}(z))=0$ (2.2.9)
which appears to imply that
$\Lambda(z)\sqrt{|p^{(0)}(z)|}=\rm{c}$ (2.2.10)
where $c$ is a constant.
It is tempting to conclude that, if $p^{(0)}(z)$ does not have any zero, then we can take $\Lambda(z)={{c}\over\sqrt{|p^{(0)}(z)}|}$. In this case the isotropy group has one element $\int{dz\over z}\Lambda(z)J(z)$ and so is an Abelian group of dimension one.
For the case of the massive particle $p^{(0)}{(z)}=1$, as a result $\Lambda=c$, and so the isotropy group contains the one generator $J_{0}$ in agreement with what we found in section (2.1). In the massless case $p^{(0)}(z)=-{1\over\sqrt{2}}(1-(z+z^{-1}))$, which also has no zeros and so one might conclude that there is an Abelian isotropy group of dimensions one. However, above we found that the isotropy group has two elements which is not in agreement with this argument.
To resolve this discrepancy we note that
@@
The problem is that this is not a convergent series as the constant term goes as $2^{2n}$ for large $n$. As a result one cannot take the square root as in equation (2.2.10) and find that the isotropy group must involve only one generator. This was the result of reference [18]. However, the isotropy group can depend on the topology that one takes for the super momenta and it is very likely that the topology we have assumed is different to that taken in reference [18], so resolving the discrepancy.
We note that equation (2.2.8) does not involve any divergent expressions and so can be used to find the isotropy group. The $\Lambda(z)$ corresponding to the isotropy group of equation (2.2.5) is given by
@@
One can verify that this does indeed satisfy equation (2.2.9) at lowest orders when taking $p^{(0)}(z)=-{1\over\sqrt{2}}(1-(z+z^{-1}))$. This problem is generic for any rest frame momenta which, when written as $p^{(0)}(z)$, possess positive and negative powers of $z$.
@@
The momenta of the general state of equation (2.2.14) are given by
@@
There is, however, a problem with the above isotropy group ${\cal H}^{3}_{m=0}$. Unlike the other isotropy groups that we will find in the rest of this paper, the one for the massless $BMS_{3}$ irreducible representation involves an infinite number of generators, and computing the commutator of the two generators gives a divergent result. One could take one of the generators to vanish on the rest frame state leaving the other to have a non-trivial action. In particular one could take the latter generator to be the one with $a=b$ which agrees with the isotropy group for the massless Poincaré group when restricted to this case. It could be that this step is required by taking the super momenta to satisfy certain topologies.
@@ 6. Summary and Discussion
$\bullet$ The massive irreducible representation of $BMS_{3}$ has the isotropy group SO(2). This agrees with that of the massive Poincaré particle in three dimensions and it contains this representation. While if we included a central charge with a very precise value the isotropy group was enhanced to SO(3). These isotropy groups agree with those found in references [18,19,20]
$\bullet$ The massless irreducible representation of $BMS_{3}$ has an isotropy group that has two generators while that for the massless Poincaré particle in two dimensions has only one. However, the commutator of the two generators is ill defined and so it is difficult to make a comparison with the massless Poincaré particle.
@@ Acknowledgement
We would like to thank Glenn Barnich for discussions and, in particular, on the role of topology of the super momenta in determining the isotropy group. RR is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101104845 (UniFlatHolo), hosted at Harvard University and Ecole Polytechnique. P.W would like to thank STFC research council for support over many years.

### v2

@@ 1 Introduction
The works of McCarthy [13, 14,15,16,17] provided a mathematical classification of the unitary irreducible representations of the BMS group in four dimensions in the absence of the super rotations. A similar analysis was performed for the BMS group in three dimensions [18,19,20] including the super rotations. The irreducible representations of $BMS_{3}$ were also discussed in reference [46] from the viewpoint of ultra relativistic limit of SL(2,R) modules. More recently the results of McCarthy were further discussed [21,22] in relation to the infrared structure of the S-matrix. Despite these results, this mathematical machinery has not yet found concrete applications in the context of scattering theory and the role of super rotations in the representation theory remains unexplored.
@@ 2 Irreducible representations of $BMS_{3}$
$J_{m},J_{n}]=i(m-n)J_{m+n}+i{Z_{1}\over 12}m(m^{2}-1)\delta_{n+m,0},\quad$
$J_{m},P_{n}]=i(m-n)P_{m+n}+i{Z_{2}\over 12}m(m^{2}-1)\delta_{n+m,0},\quad[P_{m},P_{n}]=0$ (2.0.1)
where $J_{n}$ and $P_{n}$, $n=0,\pm 1,\pm 2,\ldots$ are the super translations and the super rotations, respectively and $Z_{1}$ and $Z_{2}$ are the possible central generators. All the generators in the above commutators are related by a minus sign to the commutators in equation (5.7) given in reference [18]. The values of the central charges for three dimensional gravity in asymptotically flat spacetimes is given by $c_{1}=0$ and $c_{2}={3\over G}=-\langle Z_{2}\rangle$
@@
$\delta p_{n}=-i\sum_{m}\Lambda_{-m}(-n+m)p_{n+m}-{ic_{2}\over 12}n(n^{2}-1)\Lambda_{n}$ (2.0.5)
@@ 2.1 Massive $BMS_{3}$ representations
We will refer to this representation, which is based on the above choice of rest frame momentum, as the massive irreducible representation of $BMS_{3}$. Similarly in the next section we will refer to the massless $BMS_{3}$ irreducible representations as the one that corresponds to the same choice of momenta as for the massless irreducible representation of the Poincaré group.
@@
We recall that $\varphi_{0}=0$ and in deriving this equation we have combined the exponentials in equation (2.1.3) and the $\varphi_{n}$ in the last equation are the new variables. This irreducible representation is in the list given in [18].
@@
where we have used equation (2.1.4) and $\ldots$ means terms of order $P^{4}$ and higher. At very lowest order $p_{0}=m+\ldots$ and so at lowest order $C(p)=-p_{0}^{2}+\sum_{n,n\not=0}p_{n}p_{-n}+m^{2}=0$. Thus the super momenta of the states obey the constraint $C(p_{n})=0$ and hence the wavefunction in momentum space depends on an infinite number of variables. We note that the above general states of equation (2.1.3) do not include those with only the usual massless momentum of the Poincaré group, that is, they cannot have ${\bf p}^{\mu}{\bf p}_{\mu}=0$ with all higher super momenta being zero. This justifies our division of irreducible representations into massive and massless.
@@
Thus $X(z)$ transforms under diffeomorphisms of $z$ as a weighted scalar which is consistent with interpreting $z$ as a parameter labelling a one dimensional object. We can then write the wavefunction in x-space as $\Psi(X(z),a)$. It follows that the massive irreducible representation of $BMS_{3}$ should be viewed as an object of dimension one, that is, a string which is parameterised by $z$. We will discuss this sting interpretation in greater detail in section (3.2).
We now find the massive irreducible representation when the central charge $c_{2}\not=0$. In this case the momenta vary under the super rotations as given in equation (2.0.5). Taking into account that the rest frame momentum is given by $p_{n}^{(0)}=m\delta_{n,0}$ we find that the condition to preserve this choice of momentum is given by
${\delta p_{n}\over m}=ni(2-{c_{2}\over 12}(n^{2}-1))\Lambda_{n}=0$ (2.1.15)
In the generic case, the isotropy group is the same as above, namely the isotropy group just contains the generator $J_{0}$. However, if the central charge is such that $n=\sqrt{1+24/c_{2}}$ is an integer, then the isotropy group is ${\cal H}_{mass}^{c}=\{J_{0},J_{n},J_{-n}\}$ which is the algebra of SL(2,R). These irreducible representations are also in the list of [18]. There are only a limited number of solutions if $c_{2}$ is positive as it should be. We note that if $c_{2}$ is negative then there is only one solution, that is, $c_{2}=-24$ with $n=0$ but this does not lead to a larger isotropy group. The situation is a bit like that for supersymmetry where for certain values of the central charge the irreducible representations of supersymmetry are smaller.
If the central charge does not have one of these special values the general momentum state is given by
$|p_{n},a\rangle\equiv e^{\sum_{|n|\geq 1}\varphi_{-n}J_{n}}|p_{n}^{(0)},a\rangle$ (2.1.16)
While if $n=\pm\sqrt{1+24/c_{2}}$ is an integer then we have the above state but must set $\varphi_{n}=0=\varphi_{-n}$.
The momenta of this state are given in terms of the $\varphi_{n}$ by
${p_{n}\over m}=\delta_{n,0}+2in\varphi_{n}(1-{c_{2}\over 24}(n^{2}-1))+\sum_{p}(p^{2}-n^{2})\varphi_{-p}\varphi_{n+p}((1-{c_{2}\over 24}((n+p)^{2}-1))+\ldots$ (2.1.17)
We note that $p_{0}^{2}+\sum_{n}p_{n}p_{-n}+\dots=0$ where $\ldots$ indicates higher terms in powers of $\varphi_{n}$ and the central charge.
To construct the wavefunction in x space we must introduce a coordinate for each independent momenta and so we introduce $x_{n}$, $n\geq 1$ as well as a coordinate $y$ corresponding to the presence of the central charge generator $Z_{2}$. The transformation of these coordinates follows from demanding that
$\sum_{n}x_{-n}^{\prime}P_{n}+y^{\prime}Z_{2}=e^{-\sum_{m}\Lambda_{-m}J_{m}}(\sum_{n}x_{-n}P_{n}+yZ_{2})e^{\sum_{m}\Lambda_{-m}J_{m}}$ (2.1.18)
We find that
$\delta x_{n}=-i\sum_{m}(2m+n)\Lambda_{-m}x_{n+m},\ \ \delta y=i\sum_{n}x_{n}\Lambda_{n}n(n^{2}-1)$ (2.1.19)
We can rewrite this in terms of the quantities of equation (2.1.14) as
$\delta X(z)=-i\Lambda(z)z{dX(z)\over dz}+iz{d\Lambda(z)\over dz}X(z),\ \ \delta y=i\int dzz\Lambda(z){d^{3}\over dz^{3}}(zX(z))$ (2.1.20)
The wavefunction in x space then has the form $\Psi(X(z),y,a)$. The dependence on the coordinate $y$ could be rather trivial and it may be related to the central charge the string may carry.
@@ 2.2 Massless extended $BMS_{3}$ representations
$\delta p_{n}^{(0)}=0=-{i\over\sqrt{2}}(2n\Lambda_{n}(1+{\sqrt{2}c_{2}\over 24}(n^{2}-1))+(1-2n)\Lambda_{n-1}-(1+2n)\Lambda_{1+n})$ (2.2.2)
@@
$\quad 4\Lambda_{2}(1+{\sqrt{2}c_{2}\over 8})-3\Lambda_{1}-5\Lambda_{3}=0,\quad\quad 4\Lambda_{-2}(1+{\sqrt{2}c_{2}\over 8})-3\Lambda_{-1}-5\Lambda_{-3}=0,\ldots$ (2.2.3)
Taking $n\to-n$ in equation (2.2.2) it is easy to realise that $\Lambda_{n}=\Lambda_{-n}$. In the BMS3 case we find that equations (2.2.3) can be solved in terms of two real parameters $a$ and $b$ as follows
$\Lambda_{1}=a=\Lambda_{-1},\quad\Lambda_{0}=2b,\quad\Lambda_{2}={2\over 3}(a-b)=\Lambda_{-2},\quad\Lambda_{3}=-{(a+8b-\sqrt{2}c_{2}(a-b))\over 15}=\Lambda_{-3},\quad\ldots$ (2.2.4)
As a result the istropy group for the massless $BMS_{3}$ is ${\cal H}_{m=0}^{3}=\{H_{1},H_{2}\}$ where
$H_{1}=J_{0}-{1\over 3}(J_{2}+J_{-2})-{4\over 15}(1+{\sqrt{2}c_{2}\over 8})(J_{3}+J_{-3})+\ldots,\quad$
$H_{2}=(J_{1}+J_{-1})+{2\over 3}(J_{2}+J_{-2})-{1\over 15}(1-\sqrt{2}c_{2})(J_{3}+J_{-3})+\ldots$ (2.2.5)
As such the isotropy group has two elements and this conclusion appears to be independent of whether $c_{2}$ is zero or not.
@@
$\delta p(z)=iz\Big(\Lambda(z){d\over dz}p(z)+2p(z){d\over dz}\Lambda(z)-{c_{2}\over 12}{d^{3}\over dz^{3}}(z\Lambda(z))\Big)$ (2.2.7)
@@
$\delta p(z)=0=iz\Big(\Lambda(z){d\over dz}p^{(0)}(z)+2p^{(0)}(z){d\over dz}\Lambda(z)-{c_{2}\over 12}{d^{3}\over dz^{3}}(z\Lambda(z))\Big)$ (2.2.8)
@@
${d\over dz}(\Lambda(z)^{2}p^{(0)}(z)-{c_{2}\over 12}{d^{2}\over dz^{2}}(z\Lambda(z))=0$ (2.2.9)
As we now illustrate one has to handle this equation with care. For simplicity we will illustrate this for $c_{2}=0$. In this case it might appear that one can rewrite equation (2.2.9) as
$\Lambda(z)^{2}p^{(0)}(z)=\rm{e^{2}}\ \ {\rm or\ }\ \ \Lambda(z)\sqrt{|p^{(0)}(z)|}=\pm\rm{e}$ (2.2.10)
where $e$ is a constant.
It is tempting to conclude that, if $p^{(0)}(z)$ does not have any zeros, then we can take $\Lambda(z)={{e}\over\sqrt{|p^{(0)}(z)}|}$. In this case the isotropy group has one element $\int{dz\over z}\Lambda(z)J(z)$ and so is a group of dimension one. This is the case of the massive particle as $p^{(0)}{(z)}=1$, and as a result $\Lambda=\pm e$, and so the isotropy group contains the one generator $J_{0}$ in agreement with what we found in section (2.1). In the massless case $p^{(0)}(z)=-{1\over\sqrt{2}}(1-(z+z^{-1}))$ which has two zeros as $z={1\over 2}(1\pm\sqrt{3}i)$. Taking this value in equation (2.2.10), we find that $e=0$ and so $\Lambda=0$. This implies that the isotropy group has no elements which is not in agreement with the above two dimensional isotropy group.
We note that
@@
The problem is that this is not a convergent series as the constant term goes as $2^{2n}$ for large $n$. As a result one cannot take the square root as in equation (2.2.10). This problem is generic for any rest frame momenta which, when written as $p^{(0)}(z)$, possess positive and negative powers of $z$.
The $\Lambda(z)$ corresponding to the isotropy group of equation (2.2.5) is given by
@@
One can verify that this does satisfy equation (2.2.8) at lowest orders when taking $p^{(0)}(z)=-{1\over\sqrt{2}}(1-(z+z^{-1}))$. In fact the same divergence problem occur when we take $(\Lambda(z))^{2}$ and so one cannot reliably use equation (2.2.9), or indeed the first equation in (2.2.10). We note that equation (2.2.8) does not involve any divergent expressions and so can be used to find the isotropy group.
However, the isotropy group can depend on the topology that one takes for the super momenta and it is likely that the topology we have assumed is different to that taken in reference [18], so resolving the discrepancy.
@@
The momenta of the general state of equation (2.2.14) are given, when $c_{2}=0$, by
@@
There is, however, a problem with the above isotropy group ${\cal H}^{3}_{m=0}$. Unlike the other isotropy groups that we will find in the rest of this paper, the one for the massless $BMS_{3}$ irreducible representation involves an infinite number of generators, and computing the commutator of the two generators appears to give a divergent result. One could take one of the generators to vanish on the rest frame state leaving the other to have a non-trivial action. In particular one could take the latter generator to be the one with $a=b$ which agrees with the isotropy group for the massless Poincaré group when restricted to this case. It could be that this step is required by taking the super momenta to satisfy certain topologies.
@@ 6. Summary and Discussion
$\bullet$ The massive irreducible representation of $BMS_{3}$ has the isotropy group SO(2). This agrees with that of the massive Poincaré particle in three dimensions and it contains this representation. While if we included a central charge with a very precise value the isotropy group was enhanced to SL(2,R). These isotropy groups agree with those found in references [18,19,20]
$\bullet$ The massless irreducible representation of $BMS_{3}$ has an isotropy group that has two generators even if the central charge $c_{2}$ is non-zero. This contrasts with that for the massless Poincaré particle in three dimensions that has an isotropy generators with only one generator. However, the commutator of the two generators is ill defined and so it is difficult to make a comparison with the massless Poincaré particle. To properly understand this result requires further study.
@@
It would be interesting to find the dynamics that this string obeys. The dynamics of this string is $BMS_{3}$ invariant and so it must be invariant under the reparameterisation of the world line of equation (2.1.20) which is that for an object with an upper world index. In a Hamiltonian approach one would expect that such a local symmetry would lead to first class constrains and it would be interesting to see how these are realised on the string states.
@@ Acknowledgement
We would like to thank Glenn Barnich for discussions and, in particular, on the role of topology of the super momenta in determining the isotropy group. We also thank Blagoje Oblak for a number of very helpful comments on the first version of this paper and in particular for drawing our attention to the role played by the central charge in representation theory in three dimensions. RR is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101104845 (UniFlatHolo), hosted at Harvard University and Ecole Polytechnique. P.W would like to thank STFC research council for support over many years.